SHOW COMPLETE SOLUTION




1. water flows into a conical vessel 18cm deep and 10cm in a diameter. if the rate at which the water surface is rising is 27.52mm/s, how fast is the water flowing into the conical vesel in m^3/s when the initial depth of water is 12cm?

1
Expert's answer
2021-12-28T15:17:01-0500

1.

Let "V" denote the volume of the water in the tank.

Let "h" denote the height of the water in the tank.

Let "r" denote the radius of the water in the tank.

We are given that "V=\dfrac{1}{3}\pi r^2 h"

Since corresponding parts of similar triangles are proportional, then


"\dfrac{r}{h}=\dfrac{10/2}{18}=>r=\dfrac{5}{18}h"

"V=\dfrac{1}{3}\pi (\dfrac{5}{18}h)^2 h"

"V=\dfrac{25}{972}\pi h^3"

Differentiate both sides with respect to "t"

"\dfrac{dV}{dt}=\dfrac{25}{324}\pi h^2\dfrac{dh}{dt}"

Given "\dfrac{dh}{dt}=0.02752\ m/s, h=0.12\ m"


"\dfrac{dV}{dt}=\dfrac{25}{324}\pi (0.12)^2(0.02752)\approx9.6063\times10^{-5}\ m^3/s"